OpenStudy (anonymous):

Solve the following in [0,2pi) sin^3theta+2sin^2theta=sintheta+2

OpenStudy (anonymous):

\[\sin ^{3}\theta+2\sin ^{2}\theta=\sin \theta+2\]

OpenStudy (anonymous):

Pretend that \[\sin{\theta} = x\] solve the cubic and then resub in to solve for theta.

OpenStudy (anonymous):

so would i factor it out?

OpenStudy (anonymous):

sinx(sinx^2+2Sinx)=sinx+2

OpenStudy (anonymous):

They are saying let x = sin theta So, x^3 + 2x^2 = x + 2 Solve for x, then set that x value equal to sin theta

OpenStudy (anonymous):

ok so would i subtract x from the right side and put it on the left having the equation =2

OpenStudy (anonymous):